1. Suppose a refrigerator that sells for $700 costs $85 a year for electricity. Write an expression for the cost to buy and run this refrigerator for x years.
2. Suppose another refrigerator costs $1000 and $25 a year for electricity. Write an expression for the total cost for the refrigerator over x years.
3. Over 10 years which refrigerator costs the most? By how much?
4. In how many years will the total costs for the two refrigerators be equal?

Answer by mananth(16949)

first refrigerator cost
C(x1)=700+85x
C(10)= 700+850
C(10)= 1550
refrigerator II
C(x2) = 1000+25x
C10) =1000+250
C(10)=1250
Now you know it.
same cost
700+85x=1000+25x
60x=300
x= 5 years